OpenStudy (anonymous):

Anyone have ideas on how to solve for the angle know(-b^2+2jb)/(1-b^2+2jb) with 'j' being the complex number

OpenStudy (anonymous):

I get that \[\frac{ -b^2+2jb }{ 1-b^2+2jb }\]

OpenStudy (anonymous):

\[ \Large \frac{b(b-2i)}{(b-i)^2} \]

OpenStudy (anonymous):

Do you know how to continue from there?

OpenStudy (anonymous):

\[\Large \frac{b(b-2i)(b+i)^2}{(b-i)^2(b+i)^2} \] This will make the denominator real and you can evaluate the argument.

OpenStudy (anonymous):

if the bottom is real and the top isn't

OpenStudy (anonymous):

isn't there something like tan-1(b/a) considering a+b*j

OpenStudy (anonymous):

\[ \large \tan(\phi)=\frac{y}{x}\] Where y = Im(z) and x=Re(z)

OpenStudy (anonymous):

ah okay gotcha thanks for your help

OpenStudy (anonymous):

No Problem